Suppose you obtain a 25-year mortgage loan of $198,000 at an annual interest rate of 8.6%. The annual property tax bill is $965 and the annual fire insurance premium is $489. Find the total monthly payment for the mortgage, property tax, and fire insurance. (Round your answer to the nearest cent.)
Answer #1

L = 198,000

r = 8.6% = 0.086

t = 25 years

n=12 for monthly payment

71

198000 | 12-at 01부) 1225 1 1225

0.007166 (1 0.007166)300 1 0.007166)1 pmt 198000 300

0.007166 (1.007166)300 (1.007166)1 pmt - 198000 300

00 5102 0.007166 8.51929 nt = 198000 |-8.51929-1

0.06105 198000 751929 pmt=

pmt-198000 0.00812

prnt 1607.71..................monthly mortgage payment

.

.

annual property tax bill is $965 and the annual fire insurance premium is $489.

so total value is 965+489 = 1454

.

to find monthly value dive by 12

1454/12 = 121.166

.

so total monthly payment is

PMT = 1607.71 + 121.166

PMT 1728.88.............. total monthly payment for the mortgage